When the product of 0.2 and 0.3 is subtracted from the sum of 0.2 and 0.3, what is the difference?

I知 so confused on what to do

Answer by ikleyn(52814)

.
(0.2 + 0.3) - 0.2*0.3 = 0.5 - 0.06 = 0.44.